OpenStudy (anonymous):

Prove that this series converges or diverges, and compute its sum on a) : a) Sum {arctg[1/(n^2+n+1)]} b) Sum {sqrt(2n-1)ln(4n+1)/n(n+1)}

OpenStudy (turingtest):

from n=0 or from n=1 ?

OpenStudy (turingtest):

\[\sum_{n=0}^\infty\tan^{-1}(\frac1{n^2+n+1})\]?

OpenStudy (anonymous):

ups, a) from n=1 to infinity and b) from n=2 to..

OpenStudy (turingtest):

\[\sum_{n=1}^\infty\tan^{-1}(\frac1{n^2+n+1})\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

the book suggest tg(a-b)=[tg(a)-tg(b)]/[1+tg(a)tg(b)]

OpenStudy (anonymous):

would be useful for a)

OpenStudy (anonymous):

I thought cause book's suggestion to transform the general term in some way to related it with telescopic series, but I don't

OpenStudy (turingtest):

it's inverse tangent though in the problem, right?

OpenStudy (anonymous):

yep

OpenStudy (turingtest):

@ash2326 @KingGeorge @myininaya series help?

OpenStudy (anonymous):

what about if i put the general term like

OpenStudy (anonymous):

1/n(n+1)+1=(n-n+1)/n(n+1)+1 and do the changes n=tg(a) and n+1=tg(b), use the above formula for the tangent and turn into telescopic series? Just an idea, I'm lose...xd

OpenStudy (anonymous):

*lost

OpenStudy (anonymous):

It is well known that \[ \tan ^{-1}(x)+\tan ^{-1}(y)=\tan ^{-1}\left(\frac{x+y}{1-x y}\right)\\ x=-n\\ y= n+1\\ \tan ^{-1}(n+1)-\tan ^{-1}(n)=\tan ^{-1}\left(\frac{1}{n^2+n+1}\right)\\ Let c_n=\tan^{-1} (n)\\ then\\ a_n=\tan ^{-1}\left(\frac{1}{n^2+n+1}\right)= c_{n+1} - c_n\\ \] Use telescoping to get \[ s_n=c_{n+1}- c_1=c_{n+1}- \frac \pi 4 \]

OpenStudy (anonymous):

\[ \lim_{n\to \infty} c_{n+1} =\frac \pi 2\\ \lim_{n\to \infty} s_n= \frac \pi 2 - \frac \pi 4=\frac \pi 4 \]

OpenStudy (anonymous):

So the sum of the first integral is \(\frac \pi 4\)

OpenStudy (anonymous):

Did you understand it?

OpenStudy (anonymous):

yes, thanks a lot!

OpenStudy (anonymous):

and what about b) comparison test? limit test?

OpenStudy (anonymous):

b) solved, by limit test, if anybody wants a detailed explanation, just ask